OpenStudy (anonymous):

What value of q makes the statement true? –(q + 14) = –3 + (–14)

OpenStudy (anonymous):

@surdawi what?

OpenStudy (anonymous):

Why did you delete them @surdawi ?

OpenStudy (anonymous):

He was right..

OpenStudy (anonymous):

Will I explain u @emeliarocks12 :)

OpenStudy (anonymous):

ok @miyuru idk y he deleted his comments

OpenStudy (anonymous):

U can remove the brackets of by multiplying by -1...

OpenStudy (anonymous):

That's what @surdawi told u "Distribute"

OpenStudy (anonymous):

yeah but i a, stupid and i dont what that means @Miyuru

OpenStudy (anonymous):

Be cool it's easy don't get frustrated...:D I'll help u step by step...

OpenStudy (anonymous):

okay @Miyuru

OpenStudy (anonymous):

The first step is distribution or else removing the brackets Multiply both q and 14 by -1 be careful when multiplying. See that there is a minus out of the brackets SO both q and 14 will be minus..

OpenStudy (anonymous):

but how do i multiply an letter @Miyuru

OpenStudy (anonymous):

No problem just multiply it as a number.. :D

OpenStudy (anonymous):

so -q times -14

OpenStudy (anonymous):

@Miyuru

OpenStudy (anonymous):

Not -q times -14.. It's just -q-14... When u multiply -q and - 14 it will be 14q

OpenStudy (anonymous):

oh ok @Miyuru

OpenStudy (anonymous):

It's like this... -(a+b)=-a -b

OpenStudy (anonymous):

Now the other side that means the right side of = sign...:)

OpenStudy (anonymous):

ok so whats the answer @Miyuru

OpenStudy (anonymous):

Ohh wait don't jump to the answer at first lets try to solve it step by step.. Then u will be able to answer these type of question all by yourself

OpenStudy (anonymous):

okay @Miyuru

OpenStudy (anonymous):

i know the answer is -17 @Miyuru

OpenStudy (anonymous):

Then -q-14=-17... Still u haven't got the answer..

OpenStudy (anonymous):

okay @Miyuru

OpenStudy (anonymous):

Now take -14 to the other side...

OpenStudy (anonymous):

ok @Miyuru

OpenStudy (anonymous):

Can u tell me what did u get after taking the -14 to the right side..

OpenStudy (anonymous):

@emeliarocks12 did u get the answer...:)

OpenStudy (anonymous):

its 3 @Miyuru

OpenStudy (anonymous):

Yeah u are correct q=3 see now u git it All by yourself are u happy... U did a question alone with some help..